20) A man extends his knees as he lifts a box. When his knees reach a 80 degree angle, the center of mass of the man and the box is 20cm behind the center of rotation of his knee joint. The weight of the man (minus his lower legs and feet) is 160 pounds and the weight of the box is 40 pounds. How much torque would the mans left quadriceps muscle have to produce to maintain this position? (Hint: The man is standing on both legs. Would his left leg have to support the entire weight?)"
1
Expert's answer
2020-04-14T11:59:50-0400
"\tau=W_bd\cos{80}"

"\tau=(160\cdot 4.44822)(0.2)\cos{80}=24.7\ Nm"